A wall with an area of (3x2 + 4x + 1) square units has a rectangular window that has a width of (x + 2) units and a length of (x + 1) units. The wall also has a built-in shelving unit that occupies an area of (x2 + 5x + 6) square units of wall space. If the wall is to be covered with wallpaper, how much wallpaper will be required?
Answer by Boreal(15235)

Area = wallpaper suitable +(x+2)(x+1)+(x^2+5x+6)
x^2+3x+2+(x^2+5x+6)=2x^2+8x+8
Total area=3x^2+4x+1
area not covered=2x^2+8x+8
Area needed to be covered is x^2-4x-7 ANSWER
suppose x=10
area of windows is 132
area of shelving is 156
area of whole room is 341
area needed to be covered is 341-288=53
x^2-4x-7=53, when x=10, checks